OpenStudy (anonymous):

Is there a solution to: 1/2 (e^x + e^-x) = 4

jimthompson5910 (jim_thompson5910):

\[\Large \frac{1}{2}(e^x + e^{-x}) = 4\] \[\Large e^x + e^{-x} = 4*2\] \[\Large e^x + e^{-x} = 8\] \[\Large e^x + \frac{1}{e^{x}} = 8\] \[\Large e^{x}*(e^x + \frac{1}{e^{x}}) = 8e^{x}\] \[\Large e^{x}*e^x + e^{x}\frac{1}{e^{x}} = 8e^{x}\] \[\Large e^{2x} + 1 = 8e^{x}\] \[\Large (e^{x})^2 + 1 = 8e^{x}\] \[\Large z^2 + 1 = 8z\] Solve the last equation (use the quadratic formula) then plug in z = e^x and solve for x

OpenStudy (anonymous):

Awesome! Thank you!!

jimthompson5910 (jim_thompson5910):

no problem